y=2x
x+y=3
and find out the ordered pair

Answer by stanbon(75887)

y=2x
x+y=3
and find out the ordered pair
------------------------------------
Substitute for "y" and solve for "x":
x + 2x = 3
3x = 3
x = 1
Since y = 2x, y = 2*1 = 2
------------------------------
Solution: (1,2)
=================================
Cheers,
Stan H.